OpenStudy (anonymous):

Mr. and Mrs. Corey are newlyweds and want to purchase a home, but they need a down payment of 30000 dollars. If they want to buy their home in 3 years, how much should they save each month in their savings account that pays 6% per annum compounded monthly?

OpenStudy (anonymous):

\[1 st Month = P \] \[2nd Month = P(1+R) + P \] \[\text{3rd Month -$\quad $(P(1+R)+P)(1+R)+P= } P + P (1 + R) + P (1 + R){}^{\wedge}2\]\[=P\left(1+(1+R)+(1+R)^2\right. )\] \[\text{4th Month$\quad $- } P \left(1 + (1 + R) + (1 + R){}^{\wedge}2+(1+R)^3\right. )\] \[\text{Nth Month$\quad $- $\quad $} P\left( 1+ (1 + R) + (1 + R){}^{\wedge}2 + (1 + R){}^{\wedge}3+\text{...}\text{..}+(1+R)^{N-1}\right. )\] \[\frac{1 - (1 + R)^N}{1-(1+R)} = \frac{1-(1+R)^N}{-R}\] \[\text{ = P(} \frac{1-(1+R)^N}{-R} )\]

OpenStudy (anonymous):

\[30000=P\frac{1-\left(1+\frac{.06}{12}\right)^{36}}{-\frac{.06}{12}}\]

OpenStudy (anonymous):

P=762.66